Let f(x) = x/(x - 1), g(x) = (x - 1)/x, h(x) = 1/(x - 1) .
Then (gh)(x) =
a. 1
b. x
c. 1/x
d. -x
e. none of these

Answer by venugopalramana(3286)

Let f(x) = x/(x - 1),
g(x) = (x - 1)/x,
h(x) = 1/(x - 1)=Y SAY .
Then (gh)(x) = G(H(X)=G(Y)=(Y-1)/Y=1-1/Y=1-1/(1/(X-1)=1-(X-1)=1-X+1=-X
ANSWER IS D
a. 1
b. x
c. 1/x
d. -x
e. none of these